# Bridging Developer Instructions and Code Completion Through Instruction-Aware Fill-in-the-Middle Paradigm

### v1

@@ 9. Conclusion and Future Work
For future work, we have identified several promising directions. First, we plan to expand our instruction dataset to encompass a wider variety of programming languages beyond the current Python-dominated collection to assess IFIM’s cross-lingual generalizability. Second, while our LLM-generated instructions proved effective, we intend to explore alternative and ”wild” data sources, such as inline comments, developer forum discussions, and real-world code completion logs, which could provide richer and more diverse training signals. This will also necessitate the development of robust filtering techniques to handle inherent noise. Finally, our experiments were constrained by available computational resources; a crucial next step will be to apply and evaluate the IFIM methodology on significantly larger, state-of-the-art code LLMs to understand its scaling properties.

### v2

@@ 9. Conclusion and Future Work
For future work, we have identified several promising directions. First, we plan to expand our instruction dataset to encompass a wider variety of programming languages beyond the current Python-dominated collection to assess IFIM’s cross-lingual generalizability. Second, while our LLM-generated instructions proved effective, we intend to explore alternative and ”wild” data sources, such as inline comments, developer forum discussions, and real-world code completion logs, which could provide richer and more diverse training signals. This will also necessitate the development of robust filtering techniques to handle inherent noise. Finally, our experiments were constrained by available computational resources; an important next step will be to apply and evaluate the IFIM methodology on significantly larger, state-of-the-art code LLMs to understand its scaling properties.
Data-Availability Statement